OpenStudy (anonymous):

What is the slope of a line that passes through the point (−1, 1) and is parallel to a line that passes through (3, 6) and (1, −2)?

OpenStudy (helder_edwin):

two parallel lines have the same slope

OpenStudy (anonymous):

this not the correct answer but through di you can solve ur self

OpenStudy (anonymous):

y = mx + b m = 5 y = 5x +5 -7 = -10 + b 3 = b

OpenStudy (anonymous):

is that the slope ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok thanks

OpenStudy (helder_edwin):

what that guy did is just wrong

OpenStudy (helder_edwin):

i told u: two parallel lines have the same slope

OpenStudy (helder_edwin):

do u know how to find the slope between two points

OpenStudy (helder_edwin):

u have two points (3,6) and (1,-2) the slope is given by \[ \large \frac{y_2-y_1}{x_2-x_1} \]

OpenStudy (helder_edwin):

can u do this "chica portoriqueña"

OpenStudy (helder_edwin):

@puertoricangirl96 ???

OpenStudy (anonymous):

oh ok give me a sec.

OpenStudy (helder_edwin):

\[ \large \frac{6-(-2)}{3-1}= \]

OpenStudy (helder_edwin):

bueno adios chica portoriqueña. espero haberte ayudado lo suficiente. :)

OpenStudy (anonymous):

que nos ayudó mucho!!